# Application of Transformers for Nonlinear Channel Compensation in Optical Systems

### v2

@@ 4.5 A comparison to an LSTM-based equalizer
As discussed in Section 1, RNNs and specifically LSTMs are extensively studied for fiber nonlinearity equalization. Combination of CNN and LSTM (CNN-LSTM) generally shows better performance compared to using LSTMs alone despite incurring additional complexity. Also, more efficient LSTM structures have been introduced for single carrier and digital subcarrier multiplexing structures to use block-processing for equalization stage. To do this, the models are trained on one symbol and executed on longer blocks of symbols which results in reduction of the computational complexity [12, 11].
@@
Although a comprehensive comparison between the proposed Transformer model and CNN-LSTM structures is beyond the scope of this paper, we compare them through an example in order to demonstrate the capability of Transformers in our application. Here, we compare the performance versus complexity of the proposed Transformer with an optimized CNN-LSTM structure given in Section 3 of [11]. Table 6 shows the grid search regions for the hyper-parameters of CNN-LSTM model. We also use the same system model as shown in Fig. 6 for both networks with the block size of 128. Note that while the Transformer models are trained and evaluated at the same block-size, CNN-LSTMs are trained at block-size of 1 and evaluated at block-size of 128. Fig. 11 shows envelopes of the evaluated models performance versus complexity trade-off over different RMPS regions. The result for CNN-LSTM models with block-size equal to one is also depicted here to show how block-processing in LSTM can improve performance despite incurring much higher latency. Considering just the float-point real multiplication and ignoring latency, one can see that at higher complexities, the Transformer outperforms CNN-LSTM by a large margin (more than 0.3 dB). At lower complexities, CNN-LSTM models demonstrate a slight advantage, however, this amount is not consistent at all regions and the gap is very small. We hypothesize that at higher complexities, the attention mechanism is more powerful and can capture the system memory better than the gated LSTM structure. However, at lower complexities, the attention’s capacity is not large enough to capture the dependencies very well among the 128 symbols compared to the folding mechanism in LSTM.
Table 6: CNN-LSTM’s hyper-parameters and their values used for the grid search.
@@
Figure 11: Performance versus complexity for the Transformer and CNN-LSTM models. The blue and green lines show the envelopes for CNN-LSTM-NLC with block sizes of 1 and 128, respectively. The red line shows the envelope for Transformer-NLC.
4.6 Impact of Increasing Symbol Rate
In this part, we investigate the performance evolution of Transformer-NLC and DBP at different symbol rates. First transmission scenario in Section 4.3 with 16QAM modulation and a link with 40 spans of 80 km SSMF is considered. Additionally, the fixed Transformer-NLC model associated to Region 1 in Table 2 is considered for all simulation with 32, 64, and 96 GBaud transmission scenarios. Note that this model is close to the envelope of performance versus complexity trade-off at higher regions for 32 GBaud transmission and may not be an optimized model for higher baud rate scenarios.
The performance results for the given Transformer-NLC model is shown in Fig. 12. The nonlinear compensation gains for optimized DBP structures of various complexity operating at two samples-per-symbol are also depicted in these graphs. We can observe that the nonlinear compensation is generally more challenging at higher baud rates as the gains are shrinking. Moreover, one need to increase the complexity of DBP and use more steps-per-span to match the given Transformer-NLC model. This is in fact another issue with the application of DBP for high-speed optical transceivers ([26] and references therein) and these results confirm the advantage and flexibility of a data-driven ANN model with increasing symbol rates over the deterministic method of DBP.
(a) 32 GBaud
(b) 64 GBaud
(c) 96 GBaud
@@ 4.6 Impact of Increasing Symbol Rate
Figure 12: Performance of a given Transformer-NLC model at 32, 64, 96 GBaud at different launch powers. The results of DBP with various steps-per-span are also shown.
@@
5 Conclusion
In this work, by using Transformers, we introduced a new channel nonlinear compensation method based on the self-attention mechanism. We show that Transformers can be employed effectively for coherent long-haul transmissions. Efficient embedding and detailed implementation of Transformers as well as the use of a physics-informed mask were presented. Moreover, the designs were optimized for block-processing in order to meet the practical high-throughput/low-latency requirements of long-haul optical transmission. The presented designs were supported with numerical results which show that one can use the Transformer-NLCs as an efficient, flexible and parallelizable solution to compensate the fiber nonlinearity. A direction for further studies is simplifying the Transformer-NLC design. One possible approach is to modify the attention structure by designing a more efficient and kernel-based formulation of the self-attention. Furthermore, the impacts of quantization and pruning need to be studied for practical deployment through both post-training simplification and quantization-aware training.
@@ 5 Conclusion
Appendix A Attention Heat Maps
In this appendix, we show several two-dimensional heat maps of $(QK^{T})/\sqrt{d_{k}}$ (scaled multiplication of the queries and keys) with and without mask prior to the application of softmax function. Fig 12(a) illustrates $(QK^{T})/\sqrt{d_{k}}$ values averaged over a batch for a Transformer with three encoder layers and four heads. As seen in this figure, the majority of large values of $(QK^{T})/\sqrt{d_{k}}$ is around the main diagonal, which confirms that one can reduce the computational complexity by using the proposed masks. Thus, by using block masks we force some elements of attention matrix to zero which can reduce the computational complexity considerably. Fig. 12(c) shows the $(QK^{T})/\sqrt{d_{k}}$ values averaged over a batch for $\rho=2.6$. Furthermore, Fig. 12(b) and 12(d) shows another example of heat maps for a Transformer with three encoder layers and a single head.
Figure 13: Heat maps for the attention matrices: The 2D matrix $(QK^{T})/\sqrt{d_{k}}$ averaged over a batch. The spectrum from red to blue represents high to low values. (a) shows the attention matrix for a Transformer with 3 encoder layers and 4 heads without any masks. Rows show the values for 3 layers and columns show the values for 4 heads. (b) shows the attention matrix for a Transformer with 3 encoder layers and 1 head. Rows show the values for 3 layers. (c) and (d) are the attention matrices with the proposed mask of $\rho$ equal to 2.6 for the examples in (a) and (b), respectively.
Appendix B Impact of Model’s Hyper-Parameters
@@ B.1 Embedding Type
Table 7 shows the impact of three different embedding-generator module types. These results were obtained by a Transformer with the block size of 128, tap, key, hidden, and FFN hidden sizes of 64, and a window size of 7 for the window processing. As it can be seen from the table, CNN embedding generator module has a large positive impact on the model prediction capability with close RMPS values to the two MLP embedding generator modules.
@@
Table 7: Impact of the embedding type on performance for a Transformer with block size of 128, tap, key, hidden and FFN hidden sizes of 64, and a window size of 7.
@@ B.2 Block Size
For comparison, with an example we show the performance and complexity of Transformers with and without masks for the selected block sizes in Table 8. Note that the rest of the hyper-parameters are fixed in this case.
Table 8: Impact of block size on the Transformer-NLCs. Other hyper-parameters are as following: FFN hidden size = 64, hidden size = 64, key size = 64,window size = 15, number of heads = 4, tap size = 96.
@@
where $c_{1}=hd_{k}+3h$ and $N=\ell+b$. By expanding the above equation and after canceling out the common terms, we obtain
@@ B.6 Number of Heads
B.7 Number of Neighbors at the Output Layer

### v3

@@ IV-E Comparison to an LSTM-based equalizer
As discussed in Section I, RNNs and specifically LSTMs are extensively studied for fiber nonlinearity equalization. Combination of CNN and LSTM (CNN-LSTM) generally shows better performance compared to using LSTMs alone despite incurring additional complexity. Also, more efficient LSTM structures have been introduced for single carrier and digital subcarrier multiplexing structures to use block-processing for equalization stage. To do this, the models are trained on one symbol and executed on longer blocks of symbols which results in reduction of the computational complexity [12, 11].
@@
Although a comprehensive comparison between the proposed Transformer model and CNN-LSTM structures is beyond the scope of this paper, we compare them through an example in order to demonstrate the capability of Transformers for our application. We use an optimized CNN-LSTM structure according to the details given in Section 3 of [11]. Table III summarizes the grid search regions for the hyper-parameters of the CNN-LSTM model. We also use the same system model as shown in Fig. 6 for both networks with the block size of 128. Note that while the Transformer models are trained and evaluated at the same block-size, CNN-LSTMs are trained at block-size of 1 and evaluated at block-size of 128. Fig. 9 shows envelopes of the evaluated models’ performance versus complexity trade-off over different RMPS regions. The result for CNN-LSTM models with block-size equal to one is also depicted here to show how block-processing in LSTM can improve performance despite incurring much higher latency. Considering just the float-point real multiplication and ignoring latency, one can see that at higher complexities, the Transformer outperforms CNN-LSTM by a large margin (more than 0.3 dB). At lower complexities, CNN-LSTM models demonstrate a slight advantage, however, this amount is not consistent at all regions and the gap is small. We hypothesize that at higher complexities, the attention mechanism is more powerful and can capture the system memory better than the gated LSTM structure. However, at lower complexities, the attention’s capacity is not large enough to capture the dependencies very well among the 128 symbols compared to the folding mechanism in LSTM.
TABLE III: CNN-LSTM’s hyper-parameters and used for the grid search.
@@
Figure 9: The envelope of performance vs. complexity for NLCs with Transformer and CNN-LSTM models.
@@
Figure 10: Impact of symbol rate: performance of Model1 of Table IV compared with DBP with various steps-per-span at several launch powers. Figs (a), (b), and (c) show the results at 32, 64, 96 GBaud, respectively.
IV-F Comparison to DBP
It is well-known that DBP faces serious challenges for implementation in high-speed coherent transceivers including requiring higher over-sampling rate, lack of flexibility and generalization to the transmission scenario such as the need for more number of steps-per-spans with increasing symbol rate or at higher polarization mode dispersion (PMD) to maintain the performance. Also, use of multiple (inverse) fast Fourier transform (FFT) modules impacts the linear equalization data path. Furthermore, as found in [30], even the low-complexity DBP solutions are sensitive to quantization noise and require a high number of bit depth in order to improve over linear equalization. However, the robustness and effectiveness of neural networks under quantization and pruning are vastly shown in the literature as an advantage of such data-driven models [17, 18, 19, 31, 20].
Although a complete comparison between DBP and Transformer-NLCs is beyond the scope of this paper, we use two transmission scenarios in this section as examples to show the generalization advantage of Transformer-NLCs. We select two Transformer-NLC models close to the envelope of performance versus complexity regions of Fig. 7 and fix the model parameters for all transmission scenarios to compare them with DBP. The hyper-parameters for these models are shown in Table IV. Model1 is selected from the middle of performance versus complexity region while Model2 is chosen from the lower region. Both Transformer models use physic-informed masks with $\rho=2.6$. Performance results for DBP are presented at two samples-per-symbol with different steps-per-span (SpS) to benchmark the proposed Transformers-NLCs. Also, the transformer is trained in one launch power and is used in other launch powers by scaling according to (2). Furthermore, without loss of generality, we modified the Transformer’s output module in this section to generate the equalizer’s nonlinear distortion estimations for both polarizations in order to save in the overall computational complexity of Transformer-NLCs.
IV-F1 Impact of Increasing Symbol Rate
In this part, we investigate the performance of Transformer-NLC and DBP at different symbol rates. For this comparison, Model1 is considered for 32, 64, and 96 GBaud transmission scenarios. Note that this model is close to the envelope of performance versus complexity trade-off in the middle region for 32 GBaud transmission and may not be an optimized model for higher baud-rate scenarios.
The performance results for the given Transformer-NLC model is shown in Fig. 10. The nonlinear compensation gains for the optimized DBP structures of various complexities operating at two samples-per-symbol are also depicted in these graphs. We can observe that the nonlinear compensation is generally more challenging at higher baud rates as the gains are shrinking. Moreover, one need to increase the complexity of DBP and use more steps-per-span to match the given Transformer-NLC model. This is in fact a critical issue with the application of DBP for high-speed optical transceivers (as in [32] and references therein) and these results confirm the advantage and flexibility of a data-driven ANN model with increasing symbol rates over the deterministic method of DBP.
TABLE IV: Hyper-parameters of the two selected models of 16QAM setup for launch power sweep.
  Hyper-parameter Model1 Model2
tap size 32 12
hidden size 48 16
key size 16 16
number of heads 4 4
number of encoder layer 3 1
FFN’s hidden size 32 32
window size 7 7
IV-F2 Impact of Increasing PMD
Fig. 11 compares the performance of Model1 and Model2 at 64GBaud with DBP at three different PMD coefficient values, 0.05, 0.5 and 1.0 $ps/\sqrt{km}$, respectively. As depicted here, as the PMD coefficient increases, one need to employ DBP with higher steps-per-span to match the performance of both Transformer-NLC Models.
Figure 11: Impact of increasing PMD: performance of Model1 and Model2 transformers of Table IV at 64GB compared with DBP with various steps-per-span at several launch powers. Figs (a), (b), and (c) show the results at 64 GBaud for PMD coefficient of 0.05, 0.50, and 1.0 $ps/\sqrt{km}$, respectively.
Finally, to show a measure of complexity, Fig. 12 shows the RMPS of Model1, Model2 and DBP with different steps-per-span. The complexity of DPB per symbol is computed as
$C_{\text{DBP}}=4N_{span}N_{step}N_{up}\left(2\left(log_{2}(N_{FFT})+1\right)+1\right)$ (9)
where $N_{span},N_{step},N_{up},\text{and }N_{FFT}$ denote the number of spans, the number of steps-per-span, the oversampling ratio, and the size of FFT (which is 1024 in this work), respectively.
Figure 12: Complexity of Transformer-NLCs Model1 and Model2, and DBP with different steps-per-span
As Fig. 12 shows, the complexity of Model1 matches DBP with 3 steps-per-span while Model2 has the lowest complexity compared to any DBPs.
@@ IV-F2 Impact of Increasing PMD
V Conclusion
In this work, we introduced a new nonlinear channel compensation method based on the self-attention mechanism in Transformers. We show that Transformer-NLCs can be employed effectively for coherent long-haul transmissions. Efficient embedding and detailed implementation of Transformers as well as the use of a physics-informed mask were presented. Moreover, the designs were optimized for block-processing in order to meet the practical high-throughput/low-latency requirements of long-haul optical transmission. The presented designs were supported with numerical results for various transmission scenarios which show that one can use the Transformer-NLCs as an efficient, flexible, universal and parallelizable solution to compensate the fiber nonlinearity. To argue the benefits of Transformer-NLC, we also compared its performance versus complexity with DBP and LSTM-NLC. Specifically, through various changes in symbol rate and PMD coefficients, the advantage and flexibility of Transformer-NLC over DBP were demonstrated. A direction for further studies can be simplifying the Transformer-NLC design. One possible approach is to modify the attention structure by designing a more efficient and kernel-based formulation of the self-attention. Furthermore, the impacts of quantization and pruning need to be studied for practical deployment through both post-training simplification and quantization-aware training.
@@ V Conclusion
I Attention Heat Maps
Here, we show several two-dimensional heat maps of $(QK^{T})/\sqrt{d_{k}}$ (scaled multiplication of the queries and keys) with and without mask prior to the application of softmax function. Fig 13(a) illustrates $(QK^{T})/\sqrt{d_{k}}$ values averaged over a batch for a Transformer with three encoder layers and four heads. As seen in here, the majority of large values of $(QK^{T})/\sqrt{d_{k}}$ is around the main diagonal, which confirms that one can reduce the computational complexity by using the proposed masks. Thus, by using block masks we force some elements of attention matrix to zero which can reduce the computational complexity considerably. Fig. 13(b) shows the $(QK^{T})/\sqrt{d_{k}}$ values averaged over a batch for $\rho=2.6$. Fig. 13(c) and 13(d) shows another example of heat maps for a Transformer with three encoder layers and a single head.
Figure 13: Heat maps for the attention matrices: (a) shows the attention matrix for a Transformer with 3 encoder layers and 4 heads without any masks. Rows show the values for 3 layers and columns show the values for 4 heads. (b) shows the attention matrix for a Transformer with 3 encoder layers and one head. Rows show the values for 3 layers. (c) and (d) are the attention matrices with the proposed mask of $\rho$ equal to 2.6 for the examples in (a) and (b), respectively.
II Impact of Model’s Hyper-Parameters
@@ II-A Embedding Type
Table V shows the impact of three different embedding-generator module types. These results were obtained by a Transformer with the block size of 128, tap, key, hidden, and FFN hidden sizes of 64, and a window size of 7 for the window processing. As it can be seen from the table, CNN embedding generator module has a large positive impact on the model prediction capability with close RMPS values to the two MLP embedding generator modules.
@@
TABLE V: Impact of the embedding type on performance for a Transformer with block size of 128, tap, key, hidden and FFN hidden sizes of 64, and a window size of 7.
@@ II-B Block Size
For comparison, with an example we show the performance and complexity of Transformers with and without masks for the selected block sizes in Table VI. Note that the rest of the hyper-parameters are fixed in this case.
TABLE VI: Impact of block size on the Transformer-NLCs. Other hyper-parameters are as following: FFN hidden size = 64, hidden size = 64, key size = 64,window size = 15, number of heads = 4, tap size = 96.
@@
where $c_{1}=2hd_{k}+3h$ and $N=\ell+b$. By expanding the above equation and after canceling out the common terms, we obtain
@@ II-F Number of Heads
II-G Number of Neighbors at the Output Layer